Cos 2 76deg + Cos 2 14deg -cos76deg.cos16deg.=

Cos276deg + Cos214deg -cos76deg.cos16deg.=

Grade:11

1 Answers

Rinkoo Gupta
9 years ago
cos276+cos214-cos76cos16
(1+cos152)/2+(1+cos28)/2-cos76cos16
=1+1/2[cos(180-28)+cos28]-cos76cos16
=1+1/2[-cos28+cos28]-cos76cos16
=1-cos76cos16
=1-1/2[2cos76cos16]
=1-1/2[cos92+cos60]
=1-1/2[cos(90+20+1/2]
=1-1/2[-sin2+1/2]
=1-1/4+1/2 sin2
=3/4+(1/2) sin2
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty